# Version Age of Information with Contact Mobility in Gossip Networks

### v1

@@ III SHS Framework & Main Results
We utilize the Stochastic Hybrid Systems (SHS) framework [17] and follow the same steps as in [6] to derive the average version age $\tilde{v}_{S}$ of a subset $S$ of nodes in the network.
@@
The state at time $t$ is $\mathbf{X}(t)=[X_{0}(t),X_{1}(t),\dots,X_{n}(t)]$, where $X_{0}(t)\equiv 0$ and $X_{i}(t)$ is the version age of node $i\in\mathcal{N}$. For a transition $(i,j,k)\in\mathcal{L}$, the post‐transition state $\phi_{i,j,k}(\mathbf{X}(t))$ has the following values for its $l$th component:
$X_{l}^{\prime}=\begin{cases}X_{l}+1,&\text{if }i=0,\ j=0,\ k=\text{u},~l\in\mathcal{N},\\[6.0pt] 0,&\text{if }i=0,\ j=l,\ k=\text{u},\\[6.0pt] \min\{X_{l},X_{i}\},&\text{if }j=l,\ k=\text{g},\\[6.0pt] \min\{X_{i},X_{j}\},&\text{for }l\in\{i,j\},\ k=\text{m},\\[6.0pt] X_{l},&\text{otherwise}.\end{cases}$ (6)
@@
From now on, we consider a symmetric setting in which the following holds: $\lambda_{0j}=\lambda/n,~\forall j\in\mathcal{N}$, and hence the overall source-to-nodes update rate is $\lambda$; $\lambda_{ij}=\lambda,~\forall i,j\in\mathcal{N}$; and $\lambda_{\{i,j\}}^{m}=\lambda/f(n),~\forall i,j\in\{0\}\cup\mathcal{N}$. The function $f(n)$ controls the rate of mobility as the number of nodes grow. We focus on three main choices for $f(n)$: $n$, $c\ln n$, and $c$, for some $c>0$. We simulate gossip networks to validate the theoretical results employing an event-driven simulation to model the evolution of version age in both disconnected (DC) and fully connected (FC) network topologies. The results for simulation and corresponding theoretical values calculated using recursive methods are shown in Figs. 2, 3, 4, 5 and 6. The parameter values used for the simulation are: $\lambda_{e}\in\{0.5,1,2,5\}$ with $\lambda=1$ and $c=5$.
@@ IV Age Scaling with Contact Mobility
Figure 4: Simulated average version age of a single node $\tilde{v}_{1}$ with $f(n)=5\ln n$ in DC and FC networks.
@@
Figure 6: Simulated average version age of a single node $\tilde{v}_{1}$ with $f(n)=5$ in DC and FC networks.
@@ VI-C Disconnected Network with $f(n)=c\,\ln\,n$
$\tilde{v}_{1}=O\left(\frac{(\ln n)^{2}}{n}\right).$ (54)
@@ VI-D Fully Connected Network with $f(n)=c\,\ln\,n$
$\tilde{v}_{1}=O\left(\frac{(\ln n)^{2}}{n}\right).$ (65)

### v2

@@ III SHS Framework & Main Results
We utilize the Stochastic Hybrid Systems (SHS) framework [17] and follow the same framework and steps as in [6] to derive the average version age $\tilde{v}_{S}$ of a subset $S$ of nodes in the network.
@@
The state at time $t$ is $\mathbf{X}(t)=[X_{1}(t),\dots,X_{n}(t)]$, where $X_{i}(t)$ is the version age of node $i\in\mathcal{N}$. For a transition $(i,j,k)\in\mathcal{L}$, the post‐transition state $\phi_{i,j,k}(\mathbf{X}(t))$ has the following values for its $l^{th}$ component:
$X_{l}^{\prime}=\begin{cases}X_{l}+1,&\text{if }i=0,\ j=0,\ k=\text{u},\\[6.0pt] 0,&\text{if }i=0,\ j=l,\ k=\text{u},\\[6.0pt] \min\{X_{l},X_{i}\},&\text{if }j=l,\ k=\text{g},\\[6.0pt] \min\{X_{i},X_{j}\},&\text{for }l\in\{i,j\},\ k=\text{m},\\[6.0pt] X_{l},&\text{otherwise}.\end{cases}$ (6)
@@
From now on, we consider a symmetric setting in which the following holds: $\lambda_{0j}=\lambda/n,~\forall j\in\mathcal{N}$, and hence the overall source-to-nodes update rate is $\lambda$; $\lambda_{ij}=\lambda/(n-1),~\forall i,j\in\mathcal{N}$; and $\lambda_{\{i,j\}}^{m}=\lambda/f(n),~\forall i,j\in\{0\}\cup\mathcal{N}$. The function $f(n)$ controls the rate of mobility as the number of nodes grow. We focus on three main choices for $f(n)$: $n$, $c\ln n$, and $c$, for some $c>0$. We simulate gossip networks to validate the theoretical results employing an event-driven simulation to model the evolution of version age in both disconnected (DC) and fully connected (FC) network topologies. The results for simulation and corresponding theoretical values calculated using recursive methods are shown in Figs. 2, 3, 4, 5 and 6. The parameter values used for the simulation are: $\lambda_{e}\in\{0.5,1,2,5\}$ with $\lambda=1$ and $c=5$.
@@ IV Age Scaling with Contact Mobility
Figure 4: Simulated $\tilde{v}_{1}$ with $f(n)=5\ln n$ in DC and FC networks.
@@
Figure 6: Simulated $\tilde{v}_{1}$ with $f(n)=5$ in DC and FC networks.
@@ VI-C Disconnected Network with $f(n)=c\,\ln\,n$
$\tilde{v}_{1}=\mathcal{O}\left(\frac{(\ln n)^{2}}{n}\right).$ (54)
@@ VI-D Fully Connected Network with $f(n)=c\,\ln\,n$
$\tilde{v}_{1}=\mathcal{O}\left(\frac{(\ln n)^{2}}{n}\right).$ (65)